# Early stopping for conjugate gradients in statistical inverse problems

### v1

@@ 6 Early stopping and its prediction error
$N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d).$
In case $\lvert r_{t}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, we additionally have $N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant D-1$ such that by definition $\lvert r_{t}^{\prime}(0)\rvert<\lambda(N(\lvert r_{t}^{\prime}(0)\rvert)+1)^{-2}$. In case $\lvert r_{t}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, we have $N(\lvert r_{t}^{\prime}(0)\rvert)=D$. By monotonicity of $\lambda(\cdot)^{-2}$, we conclude
$\lvert r_{t}^{\prime}(0)\rvert\wedge\lambda_{D}^{-2}\leqslant\lambda\Big{(}\big{(}\lfloor 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}$
on the event $\bigcap_{\rho>0}\Omega(\rho)$, which has probability at least $1-e^{-z}$. ∎
@@
$k\leqslant\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{k,\lambda}+4(z+\log d)\rfloor+1,\quad k=1,\dots,d.$
@@
$k\lesssim_{p,c_{A},C_{A}}(\delta^{-2}S_{k,\lambda})^{2p/(2p+1)}+(\log d)^{2p/(2p+1)}+z^{2p/(2p+1)},\quad k=1,\dots,d.$
@@
Lemma 6.5 yields with probability at least $1-e^{-z}$
$\lambda_{k}^{-2}\leqslant\lvert r_{k}^{\prime}(0)\rvert\wedge\lambda_{D}^{-2}\leqslant\lambda\Big{(}\big{(}\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{k,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}.$
By the monotonicity of $\lambda(\cdot)^{-2}$, we obtain the first claim.
Assuming additionally (PSD($p$,$c_{A}$,$C_{A}$)) with $p>0$, we have $\lvert r_{k}^{\prime}(0)\rvert\geqslant\sum_{i=1}^{k}\lambda_{i}^{-2}\gtrsim_{p,C_{A}}k^{2p+1}$ such that we can conclude the second claim again by Lemma 6.5 on the same event. ∎
@@
Consider the Gaussian noise model (GN). Let $\sigma$ be any random variable taking values in $(0,d]$ (possibly depending on $\xi$). Without any assumption on $(\lambda_{i})$, we have for all $z\geqslant 1$ with probability at least $1-2e^{-z}$
@@
for $k=1,\dots,d$, $i=0,\dots,d$. Note that each $Q_{g,k,i}$ is an at most $(k+1)$-dimensional subspace, since $\operatorname{Pol}_{k}$ has dimension $k+1$. We define
@@
$r_{\sigma,<}g=\sum_{j=i_{\sigma}+1}^{d}r_{\sigma}(\tilde{\lambda}_{j}^{2})\sum_{l=1,\dots,D:\,\lambda_{l}=\tilde{\lambda}_{j}}\langle{g},{u_{l}}\rangle u_{l}\in Q_{g,\lceil\sigma\rceil,i_{\sigma}}.$
@@
$\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert=\lVert\mathsf{r_{\sigma,<}}g\rVert\frac{\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert}{\lVert\mathsf{r_{\sigma,<}}g\rVert}\leqslant\lVert\mathsf{r_{\sigma,<}}g\rVert\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert,$ (6.5)
we can reduce the problem to bounding the norm of the projection $\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert=\delta\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}Z\rVert$. Note that $\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert]\leqslant\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert^{2}]^{1/2}\leqslant\sqrt{k+1}$ because $\lVert\Pi_{g,k,i}Z\rVert^{2}$ is $\chi^{2}$-distributed with $\dim(Q_{g,k,i})$ degrees of freedom for deterministic $k$ and $i$. Moreover, the function $x\mapsto\lVert\Pi_{g,k,i}x\rVert$, $x\in{\mathbb{R}}^{D}$, is $1$-Lipschitz such that we obtain by Gaussian concentration [18, Theorem 3.4] and Bonferroni’s inequality
[…]
$\leqslant\sum_{k=1}^{d}\sum_{i=0}^{d}\operatorname{{\mathbb{P}}}\Big{(}\lVert\Pi_{g,k,i}Z\rVert-\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert]\geqslant\big{(}2z+2\log(d(d+1))\big{)}^{1/2}\Big{)}$
$\leqslant d(d+1)e^{-(2z+2\log(d(d+1)))/2}=e^{-z}.$
@@
$\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert$ $\leqslant\delta\Big{(}\sqrt{\lceil\sigma\rceil+1}+\big{(}2z+2\log(d(d+1))\big{)}^{1/2}\Big{)}$
@@
also with probability at least $1-e^{-z}$, where we used that $t\mapsto S_{t,\lambda}$ is nondecreasing by 3.2(c). Combining (6.5)–(6.7) yields the first claim. For the second assertion, we use $\lvert r_{\lfloor\sigma\rfloor}^{\prime}(0)\rvert\leqslant\lvert r_{\sigma}^{\prime}(0)\rvert<\lambda_{D}^{-2}$ by assumption and obtain by Corollary 6.6 under (PSD($p$,$c_{A}$,$C_{A}$))
@@
$\operatorname{{\mathbb{E}}}\bigg{[}\frac{\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert^{2}}{\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}}\bigg{]}\lesssim\delta^{2}\big{(}\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}]+\log d\big{)}.$
@@
$\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]$ $\leqslant\operatorname{{\mathbb{E}}}\big{[}\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}\big{]}^{1/2}\operatorname{{\mathbb{E}}}\bigg{[}\frac{\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert^{2}}{\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}}\bigg{]}^{1/2}$
@@
with a constant $C_{\mu,p,C_{A},C_{\kappa}}>0$ depending only on $\mu$, $p$, $C_{A}$ and $C_{\kappa}$. Under (PSD($p$,$c_{A}$,$C_{A}$)), this gives the minimax optimal rate for all regularities ${\mu>0}$ such that $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
@@ 7.1 Bound for the reconstruction error
$=\lVert\widehat{f}_{t}-f^{\dagger}\rVert_{<}^{2}+\lVert\widehat{f}_{t}-f^{\dagger}\rVert_{>}^{2}$
@@
$\leqslant 2\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}+2\lVert A^{\dagger}(I_{D}-r_{t}(AA^{\top}))_{<}\xi\rVert^{2}+\lvert r_{t}^{\prime}(0)\rvert\lVert A(\widehat{f}_{t}-f)\rVert^{2},$ (7.1)
@@
$\lVert A^{\dagger}(I_{D}-r_{t}(AA^{\top}))_{<}\xi\rVert^{2}=\lVert\mathsf{x^{-1/2}(1-r_{t})_{<}}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert S_{t,\lambda}.$ (7.2)
@@
Assume the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$ for the minimum-norm solution $f^{\dagger}$. Then, for $t\in[0,d]$, we have
@@
By Proposition 7.1, it remains to bound $\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}$. Note that $\langle{f^{\dagger}},{v_{i}}\rangle=f_{i}$ for $i=1,\dots,D$ and $\langle{f^{\dagger}},{v_{i}}\rangle=0$ for $i=D+1,\dots,P$ in case of $P>D$. Using the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$ and Hölder’s inequality with the conjugate exponents $\tilde{p}\coloneqq(2\mu+1)/(2\mu)$ and $\tilde{q}\coloneqq 2\mu+1$, we obtain
@@
Under (DN) and the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, we have for $t\in(0,d]$
@@
Assume (PSD($p$,00,$C_{A}$)) with $p>1/2$ and (GN). Under the source condition $\lVert f^{\dagger}\rVert_{\mu}\leqslant R$, the reconstruction error at the balanced oracle satisfies
@@
with a constant $C_{\mu,p,C_{A}}>0$ depending only on $\mu$, $p$ and $C_{A}$. In particular, under (PSD($p$,$c_{A}$,$C_{A}$)), the balanced oracle is simultaneously minimax adaptive for prediction and reconstruction error for all $\mu>0$ with $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
@@
Corollary 7.3 and the identity $S_{\tau_{\mathfrak{b}},\lambda}=A_{\tau_{\mathfrak{b}},\lambda}$ yield
@@
where we have applied Jensen’s inequality. The first summand has the desired rate by Theorem 5.8. Using 5.3(d), (5.4) and (5.6), we deduce similarly to the proof of Theorem 5.8
@@
$\leqslant\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu}+C_{A}^{2}\delta^{2}\tfrac{2p}{2p-1}\rho^{(2p+1)/(2p)}\big{)}$
@@
The error rate $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ coincides with the minimax rate in the infinite-dimensional Gaussian sequence space model, and the corresponding lower bound can be proven in the same way [see 15, Proposition 4.23], very similar to the proof of Proposition 5.11: under (PSD($p$,$c_{A}$$\infty$)) with $p>0$ and (GN), the rate $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ is minimax optimal, provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$. If $D\ll(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$, then $\delta^{2}D^{2p+1}\ll\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$, and we obtain better error bounds without regularisation exactly as in Remark 5.12:

### v2

@@ 6 Early stopping and its prediction error
$N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d).$ (6.4)
In case $\lvert r_{t}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, we additionally have $N(\lvert r_{t}^{\prime}(0)\rvert)\leqslant D-1$ such that by definition $\lvert r_{t}^{\prime}(0)\rvert<\lambda(N(\lvert r_{t}^{\prime}(0)\rvert)+1)^{-2}$. By monotonicity of $i\mapsto\lambda(i)^{-2}$ and (6.4), we obtain
$\lvert r_{t}^{\prime}(0)\rvert<\lambda\Big{(}\big{(}\lfloor 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}.$
In case $\lvert r_{t}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, we have $N(\lvert r_{t}^{\prime}(0)\rvert)=D$ such that (6.4) yields
$D\leqslant\lfloor 2\delta^{-2}(1-e^{-1})^{-1}S_{t,\lambda}+4(z+\log d)\rfloor.$
Therefore, both asserted implications hold for all $t\in[0,d]$ on the same event $\bigcap_{\rho>0}\Omega(\rho)$, which has probability at least $1-e^{-z}$. ∎
@@
$k<\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{k,\lambda}+4(z+\log d)\rfloor+1,\quad k=0,\dots,d.$
@@
$k\lesssim_{p,c_{A},C_{A}}(\delta^{-2}S_{k,\lambda})^{2p/(2p+1)}+(\log d)^{2p/(2p+1)}+z^{2p/(2p+1)},\quad k=0,\dots,d.$
@@
Consider the event of Lemma 6.4, which holds with probability at least $1-e^{-z}$. For $k=0$, there is nothing to prove. Let $k=1,\dots,d$. In case $\lvert r_{k}^{\prime}(0)\rvert<\lambda_{D}^{-2}$, we then have, by the first implication of Lemma 6.4, the strict inequality
$\lambda_{k}^{-2}\leqslant\lvert r_{k}^{\prime}(0)\rvert<\lambda\Big{(}\big{(}\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{k,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}.$
By the monotonicity of $i\mapsto\lambda(i)^{-2}$, we obtain the first claim. In case $\lvert r_{k}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}$, noting that $k\leqslant D$, the asserted bound is directly given by the second implication of Lemma 6.4.
Assuming additionally (PSD($p$,$c_{A}$,$C_{A}$)) with $p>0$, we have $\lvert r_{k}^{\prime}(0)\rvert\geqslant\sum_{i=1}^{k}\lambda_{i}^{-2}\gtrsim_{p,C_{A}}k^{2p+1}$, $k=1,\dots,d$, such that we can conclude the second claim again by the first implication of Lemma 6.4 on the same event. ∎
@@
Consider the Gaussian noise model (GN). Let $\sigma$ be any random variable taking values in $[0,d]$ (possibly depending on $\xi$). Without any assumption on $(\lambda_{i})$, we have for all $z\geqslant 1$ with probability at least $1-2e^{-z}$
@@
for $k,i=0,\dots,d$. Note that each $Q_{g,k,i}$ is an at most $(k+1)$-dimensional subspace, since $\operatorname{Pol}_{k}$ has dimension $k+1$. Recall that we set $x_{1,0}\coloneqq\infty$ and define
@@
$\mathsf{r_{\sigma,<}}g=\sum_{j=i_{\sigma}+1}^{d}r_{\sigma}(\tilde{\lambda}_{j}^{2})\sum_{l=1,\dots,D:\,\lambda_{l}=\tilde{\lambda}_{j}}\langle{g},{u_{l}}\rangle u_{l}\in Q_{g,\lceil\sigma\rceil,i_{\sigma}}.$
@@
$\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert=\lVert\mathsf{r_{\sigma,<}}g\rVert\frac{\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert}{\lVert\mathsf{r_{\sigma,<}}g\rVert}\leqslant\lVert\mathsf{r_{\sigma,<}}g\rVert\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert$ (6.7)
for $\mathsf{r_{\sigma,<}}g\neq 0$, we can reduce the problem to bounding the norm of the projection $\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert=\delta\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}Z\rVert$. If $\mathsf{r_{\sigma,<}}g=0$, the bound (6.7) on $\lvert\langle{\xi},{\mathsf{r_{\sigma,<}}g}\rangle\rvert=0$ continues to hold. In case $\dim(Q_{g,k,i})>0$ for deterministic $k$ and $i$, note that $\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert]\leqslant\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert^{2}]^{1/2}\leqslant\sqrt{k+1}$ because $\lVert\Pi_{g,k,i}Z\rVert^{2}$ is $\chi^{2}$-distributed with $\dim(Q_{g,k,i})$ degrees of freedom. Otherwise, $\lVert\Pi_{g,k,i}Z\rVert^{2}=0$. Moreover, the function $x\mapsto\lVert\Pi_{g,k,i}x\rVert$, $x\in{\mathbb{R}}^{D}$, is $1$-Lipschitz such that we obtain by Gaussian concentration [18, Theorem 3.4] and Bonferroni’s inequality
$\operatorname{{\mathbb{P}}}\Big{(}\max_{0\leqslant k,i\leqslant d}\big{(}\lVert\Pi_{g,k,i}\xi\rVert-\delta\sqrt{k+1}\big{)}\geqslant\delta\big{(}2z+4\log(d+1)\big{)}^{1/2}\Big{)}$
$\leqslant\sum_{k,i=0}^{d}\operatorname{{\mathbb{P}}}\Big{(}\lVert\Pi_{g,k,i}Z\rVert-\operatorname{{\mathbb{E}}}[\lVert\Pi_{g,k,i}Z\rVert]\geqslant\big{(}2z+4\log(d+1)\big{)}^{1/2}\Big{)}$
$\leqslant(d+1)^{2}e^{-(2z+4\log(d+1))/2}=e^{-z}.$
@@
$\lVert\Pi_{g,\lceil\sigma\rceil,i_{\sigma}}\xi\rVert$ $\leqslant\delta\Big{(}\sqrt{\lceil\sigma\rceil+1}+\big{(}2z+4\log(d+1)\big{)}^{1/2}\Big{)}$
@@
also with probability at least $1-e^{-z}$, where we used that $t\mapsto S_{t,\lambda}$ is nondecreasing by 5.3(d). Combining (6.7)–(6.9) yields the first claim. For the second assertion, we use $\lvert r_{\lfloor\sigma\rfloor}^{\prime}(0)\rvert\leqslant\lvert r_{\sigma}^{\prime}(0)\rvert<\lambda_{D}^{-2}$ by assumption and obtain by Corollary 6.5 under (PSD($p$,$c_{A}$,$C_{A}$))
@@
$\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]\lesssim\delta\operatorname{{\mathbb{E}}}\bigg{[}\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert\Big{(}\sqrt{\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}}+\sqrt{\log d}\Big{)}\bigg{]}.$
@@
$\operatorname{{\mathbb{E}}}[\lvert\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert]$ $\lesssim\delta\operatorname{{\mathbb{E}}}\big{[}\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}\big{]}^{1/2}\operatorname{{\mathbb{E}}}[\delta^{-2}S_{\tau_{\mathfrak{b}},\lambda}+\log d]^{1/2}$
@@
with $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ given in (3.2) and a constant $C_{\mu,p,C_{A},C_{\kappa}}>0$ depending only on $\mu$, $p$, $C_{A}$ and $C_{\kappa}$. Under (PSD($p$,$c_{A}$,$C_{A}$)), this gives the minimax optimal rate for all regularities ${\mu>0}$ such that $\sqrt{D}\lesssim(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
@@ 7.1 Bound for the reconstruction error
$=\lVert\widehat{f}_{t}-f^{\dagger}\rVert_{<}^{2}+\lVert\widehat{f}_{t}-f^{\dagger}\rVert_{\geqslant}^{2}$
@@
$\leqslant 2\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}+2\lVert A^{\dagger}(I_{D}-r_{t}(AA^{\top}))\xi\rVert_{<}^{2}+\lvert r_{t}^{\prime}(0)\rvert\lVert A(\widehat{f}_{t}-f)\rVert^{2},$ (7.1)
@@
$\lVert A^{\dagger}(I_{D}-r_{t}(AA^{\top}))\xi\rVert_{<}^{2}=\lVert\mathsf{x^{-1/2}(1-r_{t})_{<}}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}\leqslant\lvert r_{t}^{\prime}(0)\rvert S_{t,\lambda}.$ (7.2)
@@
Assume the source condition $\lVert f\rVert_{\mu}\leqslant R$. Then, for $t\in[0,d]$, we have
@@
By Proposition 7.1, it remains to bound $\lVert r_{t,<}(A^{\top}A)f^{\dagger}\rVert^{2}$. Note that $\langle{f^{\dagger}},{v_{i}}\rangle=f_{i}$ for $i=1,\dots,D$ and $\langle{f^{\dagger}},{v_{i}}\rangle=0$ for $i=D+1,\dots,P$ in case of $P>D$. Using the source condition $\lVert f\rVert_{\mu}\leqslant R$ and Hölder’s inequality with the conjugate exponents $\tilde{p}\coloneqq(2\mu+1)/(2\mu)$ and $\tilde{q}\coloneqq 2\mu+1$, we obtain
@@
Under (DN) and the source condition $\lVert f\rVert_{\mu}\leqslant R$, we have for $t\in(0,d]$
@@
Assume (PSD($p$,00,$C_{A}$)) with $p>1/2$ and (GN). Under the source condition $\lVert f\rVert_{\mu}\leqslant R$, the reconstruction error at the balanced oracle satisfies
@@
with $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ given in (3.5) and a constant $C_{\mu,p,C_{A}}>0$ depending only on $\mu$, $p$ and $C_{A}$. In particular, under (PSD($p$,$c_{A}$,$C_{A}$)), the balanced oracle is simultaneously minimax adaptive for prediction and reconstruction error for all $\mu>0$ with $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$.
@@
In the case $Af=0$, we have $\tau_{\mathfrak{b}}=0$ such that $\widehat{f}_{\tau_{\mathfrak{b}}}=f^{\dagger}=0$ and the bound is clear. Suppose $Af\neq 0$. Then $\tau_{\mathfrak{b}}>0$ and $\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert>0$. Corollary 7.3 and the identity $S_{\tau_{\mathfrak{b}},\lambda}=A_{\tau_{\mathfrak{b}},\lambda}$ yield
@@
where we have applied Jensen’s inequality. The first summand has the desired rate by Theorem 5.8. Using 5.3(c), (5.4) and (5.6), we deduce similarly to the proof of Theorem 5.8
@@
$\leqslant\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu}+(C_{A}\vee 1)^{2}\delta^{2}\tfrac{2p}{2p-1}\rho^{(2p+1)/(2p)}\big{)}$
@@
The error rate $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ coincides with the minimax rate in the infinite-dimensional Gaussian sequence space model, and the corresponding lower bound can be proven in the same way [see 15, Proposition 4.23], very similar to the proof of Proposition 5.11: under (PSD($p$,$c_{A}$$\infty$)) with $p>0$ and (GN), the rate $\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$ is minimax optimal over signals $f$ satisfying $\lVert f\rVert_{\mu}\leqslant R$, provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$. If $D\ll(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}$, then $\delta^{2}D^{2p+1}\ll\mathcal{R}^{*,\mathrm{rec}}_{\mu,p,R}(\delta)$, and we obtain exactly as in Remark 5.12: